A 7.6 kg chair is pushed across a frictionless floor with a force of 42 N that is applied at an angle of 22° downward from the horizontal. What is the magnitude of the acceleration of the chair?
1
Expert's answer
2010-12-05T06:15:47-0500
The horizontal projection of the force applied to the chair is
Fh = Fcos (α) = 42 cos(22°) = 38.9 N.
According ot the sewcond Newton's law,
a = Fh/m = 38.9 / 7.6 = 5.11 m/s2